OpenStudy (anonymous):

f(x)=2(x-6)^2+4 how do u convert f(x) in to standard form

OpenStudy (mayankdevnani):

\[\large \bf f(x)=2(x-6)^2+4\] just open the bracket,

OpenStudy (mayankdevnani):

\[\large \bf (a-b)^2=a^2+b^2-2ab\]

OpenStudy (mayankdevnani):

so we get, \[\large \bf f(x)=2(x^2+36-12x)+4\]

OpenStudy (anonymous):

mayankdevnain could u please help me

OpenStudy (mayankdevnani):

understood? @ladystarkzombieslayer

OpenStudy (mayankdevnani):

sure.. @zaqwsxedcrfv

OpenStudy (anonymous):

come to my question pl

OpenStudy (mayankdevnani):

now,expand, \[\large \bf 2x^2+72-24x+4\] and then solve rest of the part